# Generating a Contact Matrix for Aged Care Settings in Australia: an agent-based model study

### v1

@@ 2.2.7 Activity Generation and Execution
Activity generation for both staff and resident agents is managed through specific functions invoked by the statecharts or other elements of the simulation logic. The activity generation process involves defining both the activation of the specific activity, including the prioritisation and the creation of its token. Within the agent a process model reads the generation function and locates the activity’s priority. Both staff and resident models utilise priority-based functions to determine the order of task execution. In addition, activity tokens are employed within the model to represent the discrete units of work or actions that an agent must perform. Each token encapsulates the details of a specific activity, including the required subordinate actions, the sequence in which they should occur, location of activity and any associated timing constraints. The tokens are carriers of activity-specific information, tokens ensure that the agent performs activities according to the defined parameters even between agents. The management of activity tokens involves queuing, prioritisation, and execution, with tokens queued based on their priority. Each agent’s process model selects the appropriate token for execution, guiding the agent’s actions and ensuring consistency and adherence to the activities’ requirements.
@@ 2.2.8 Social Force Model and Pedestrian Dynamics
A pedestrian simulation using the well characterised Helbing-Molnár-Farkas-Vicsek social force models is used to implement the movement of agents within the simulation environment (Helbing2000Panic). This model differs from other social force models by incorporating a more comprehensive set of interaction forces, including individual desires to reach specific destinations while avoiding collisions with other pedestrians and obstacles. It also accounts for the variability in pedestrian behaviour, making it particularly effective for simulating realistic movement patterns in complex environments (Helbing2000Panic). The pedestrian model within the simulation is triggered when an activity template requires an agent to navigate to a specific location, such as moving between rooms or transitioning to different areas. The process model within the staff or resident activity generates movement details as part of the task parameters, including the destination, path, and required speed. These parameters are passed to the pedestrian model through the activity templates, which handles the agent’s physical navigation. The pedestrian model ensures realistic movement by managing timing, sequence, and interactions with obstacles and other agents. Using a social force approach, it calculates desired speed, directs movement toward attraction points, and programs agents to follow the shortest routes while avoiding collisions. These features collectively enhance the accuracy of agent interactions and behaviours within the simulation.
@@ 2.2.9 Subcomponent Integration
The execution of tasks follows the logic defined by the process model, activity templates, and the pedestrian model. Once an activity token is selected for execution, the process model guides the agent through each step, referencing the activity template for specific instructions. The execution logic ensures that tasks are carried out in the correct order, with appropriate timing and resource management. The agent’s state is continuously updated to reflect progress, and any deviations or interruptions are managed according to predefined rules. The integration with the pedestrian model during execution allows for coordinated movement and task completion, ensuring that the agent’s actions are synchronised with their location and performed within the appropriate context.
@@ 2.2.10 Feedback Mechanisms
Figure 2: Task flow of the agents within the aged care home.
@@ 2.2.11 Integrated Model Operation
For staff, the process models are tailored to manage multiple, often concurrent, activities that may have conflicting priorities. Staff agents are frequently tasked with balancing operational duties, such as logistical support, administrative functions, and direct care activities, in addition to balancing unscheduled events [Table 3]. The process model is designed to adjust execution of tasks as they emerge through priority-based queues, ensuring that tasks are addressed in a logical order based on urgency and importance. Each task is assigned a priority level, and the process model dynamically manages the priority queue, continuously re-evaluating tasks as new ones are generated, or existing tasks are completed. This process model also allows the model to effectively handle conflicting tasks, ensuring the lower priority tasks are either delayed or requeued as necessary. This is particularly important as staff members often juggle multiple operational responsibilities that may require immediate attention. In addition, an importance threshold was assigned in the model to manage the prioritisation-based system. Tasks can be interrupted if they exceed the threshold and are higher priority than the current executing task [Table 3]. In addition, if the interrupted task is still within the importance threshold, then it is requeued back into the process model to be completed.
@@ 2.4.1 Identifying Simulated Contacts
We simulated the movement of all the agent types, capturing their interactions and proximity throughout each staff shift. This enabled us to directly compute the instances of close contact among all individuals present, along with the duration of these interactions. In addition, as multiple studies on respiratory illnesses revealed that droplets extend far beyond 1.5 metres which is the recommended social distancing policy for the World Health Organization (bourouiba2014sneeze; bahl2022precautions; xie2007droplets). As a result, separate contact analyses were performed for contact radii of 1.5 and 3 metres to analyse differences in nonphysical contacts within the space. In the model, a contact is defined as an event where two agents come within 1.5 or 3 metres of each other for a duration equal to or exceeding 3 seconds. Mathematically, a contact between two agents i and j is represented as:
@@
Within the model, the contact data record is recorded when agent type is with the re and within the same room to avoid contacts occurring through physical layout. The data structure keeps track of contacts over time. Contacts can persist through consecutive time-steps until the point where the two agents are no longer in close proximity. If the same agents come into contact again after a period of separation, a new contact record is created and logged for that pair of agents. This new contact record will have a new start timestamp. The process continues in this manner, with new contact records being generated for each instance of contact between the same pair of agents, and timestamps being updated accordingly. The contact record, once created, includes the following parameters: (i) the identifier of the contacter; (ii) the identifier of the contactee; (iii) the start time of the contact within the effective radius; (iv) the end time of the contact; and (v) the duration of the contact.
@@
$A_{i,k}=\frac{\sum_{i\in k}A_{i,l}}{N_{k}}$
where $k$ denotes agent types and $N_{k}$ is the number of agents of type $k$.
@@ 2.5 Transmission Risk Simulation
To complement the construction of contact matrices and provide a framework for interpreting their epidemiological implications, a probabilistic airborne transmission model was integrated into the simulation to calculate infection risk based on agent proximity and cumulative contact duration. Transmission probability was dynamically evaluated for each susceptible-infectious agent pair during the simulation based on recorded spatial proximity and accumulated exposure time. Infection outcomes were determined based on the calculated transmission probability at the end of each recorded contact event.
@@ 2.5.1 Initial Seeding
To initiate the transmission simulations within the model, an initial infection seeding strategy was employed. For each simulation run, a single staff agent was randomly selected at the beginning of a designated shift to serve as the primary infection case. Only one initial infection was seeded per 24-hour simulation cycle. Separate simulation sets were conducted independently for each shift. Each index case was assumed to be infectious from the start of their respective shift, and all infection events were recorded across the 24-hour simulation period. To allow for direct comparison between scenarios, identical random seeds were used across intervention runs, ensuring that all shared stochastic processes, such as agent movement, task allocation, and seeding, remained consistent, isolating the effect of the intervention.
@@ 2.5.2 Transmission Dynamics
A probabilistic transmission risk model, applying a proximity-based Wells-Riley formulation, was integrated into the simulation (riley1978measles). Transmission risk between agents was calculated based on direct proximity and cumulative contact duration, without assuming homogeneous aerosol mixing or dilution effects through room-scale ventilation. The transmission probability P(i,j) from an infectious individual i to a susceptible individual j was defined as:
@@
Where $\lambda$ is the effective transmission rate per second of close contact, calculated as the product of the quanta generation rate (I) and the inhalation rate (q) and normalised to a per-second basis; t(i,j) represents the cumulative contact duration, measured in seconds, between individuals i and j, derived directly from simulation-recorded interaction data [Table 5]
@@
I Quanta generation rate 25 quanta/hr (95% CI: 14–54) (buonanno2020quantitative; morawska2020airborne; dai2020ventilation)
q Inhalation rate of susceptible individuals during sedentary activities 0.5 m3/hr (95% CI: 0.48–0.54) (buonanno2020quantitative; guo2020aerosol; usepa2011handbook)
@@ 2.5.4 Vaccination Intervention Scenarios
Given the elevated vulnerability of aged care settings and the recognised role of booster vaccination in mitigating infection risk, intervention scenarios incorporating resident and staff booster coverage were evaluated. Vaccination status was incorporated into the simulation as an agent attribute assigned at initialisation. For each simulation run, staff agents were assigned a binary vaccination status according to the intervention scenario being evaluated. Vaccination status remained fixed throughout the 24-hour simulation cycle.
During the evaluation of transmission probabilities, vaccination status was incorporated as a scaling adjustment applied to the calculated probability of infection for susceptible agents. If a susceptible agent was recorded as vaccinated, their computed transmission probability Ptransmission(i,j) was adjusted as:
@@
Booster vaccination was modelled assuming receipt of the Pfizer-BioNTech Comirnaty Omicron XBB.1.5 formulation. This vaccine was selected as it represented the predominant COVID-19 booster administered in Australia during the study period, consistent with Australian Technical Advisory Group on Immunisation recommendations prioritising mRNA-based vaccines for booster doses and confirmed in national policy documents and vaccine rollout reports (atagi2025booster; ausgov2025guidance; tga2025approval). Vaccine effectiveness against any SARS-CoV-2 infection for the Comirnaty Omicron XBB.1.5 booster was assumed to be 70.6% (95% CI: 43.3%–84.8%), based on recently reported effectiveness estimates encompassing both symptomatic and asymptomatic infections (ward2025comirnaty). All adjustments to infection probability based on vaccination status were applied dynamically during the simulation. Vaccination status modified the probability of infection at the time of contact evaluation without altering other agent attributes, including mobility patterns, task execution, or contact behaviours. Within the simulation, vaccination status modified transmission probability only at the point of contact event evaluation. At the conclusion of each cumulative contact between an infectious and susceptible agent, the adjusted transmission probability $P_{\text{transmission}}^{*}(i,j)$ was computed based on the susceptible agent’s vaccination status. A uniform random number $U\sim\mathcal{U}(0,1)$ was then drawn to determine infection status, following the method described above. Vaccination status did not modify other agent attributes, including movement patterns, task execution, contact rates, or behaviour. All adjustments were applied dynamically, ensuring consistency of simulation logic across vaccinated and unvaccinated agents except for the reduced risk of infection upon exposure. Three vaccination intervention scenarios were evaluated: (i) no booster vaccination among staff or residents (baseline); (ii) booster vaccination administered to all residents but not staff; and (iii) booster vaccination administered to both staff and residents. Vaccination status was assigned at model initialisation and fixed throughout the 24-hour simulation cycle. All infection probability adjustments were applied based on vaccination status as previously described.
@@ 3.2 Contact Matrix
For the contact matrix generation, the simulation ran for 5 days, starting at 6:15 AM on day 1 and ending at 6:15 AM on day 6. During this period, multiple data series were collected for the analysis, specifically focusing on two different contact radii: 1.5 meters and 3 meters [Figure 3]. At a 1.5-meter contact radius, simulated interactions among residents showed variation across different care levels. Residents in medium care had the highest predicted intra-group interactions, averaging 64.4 contacts [95% UI: 52.5–80.1], followed closely by low care residents with 49.6 contacts [95% UI: 41.7–56.3]. High care residents had substantially fewer intra-group contacts, averaging 3.5 [95% UI: 2.3–5.0]. Staff-resident interactions were most frequent for shift 1 and shift 2 care workers, particularly with medium care residents (60.9 [95% UI: 45.0–73.3] and 58.1 [95% UI: 33.3–72.0] contacts, respectively). Contacts with low care residents were also high for shift 1 (54.3 [95% UI: 41.3–61.9]) and shift 2 (52.0 [95% UI: 28.2–64.6]). In contrast, shift 3 staff demonstrated the lowest levels of interaction across all resident types, with predicted contacts not exceeding 35 across any group and intra-shift contacts averaging only 2.5 [95% UI: 0.6–6.4].
@@
At a 3-meter contact radius, the simulated interaction patterns largely mirrored those observed at 1.5 meters, though overall contact frequencies were slightly increased in most groups. Medium care residents had the highest intra-group interactions, averaging 64.8 contacts [95% UI: 45.3–69.1], followed closely by low care residents with 51.0 contacts [95% UI: 47.8–52.1]. High care residents continued to exhibit fewer intra-group contacts, averaging 3.0 [95% UI: 2.5–4.6]. Among staff, shift 1 had the most frequent predicted interactions, particularly with medium care residents (72.9 [95% UI: 49.3–81.2]) and low care residents (66.1 [95% UI: 44.8–72.1]). Shift 2 staff showed similar patterns, with 60.0 contacts with medium care residents [95% UI: 49.9–65.2] and 54.4 with low care residents [95% UI: 52.5–59.1]. Shift 3 staff maintained the lowest levels of interaction across all groups, with resident contacts not exceeding 36 and intra-shift contacts averaging only 4.7 [95% UI: 4.2–6.0].
@@
For the 1.5-meter contact radius, the GLM analysis applied to the simulated contact data revealed that low care residents had a significantly greater likelihood of being involved in interactions compared to high care residents, with a coefficient of 1.21 [95% CI: 0.97 to 1.37]. Medium care residents also showed elevated predicted interaction frequencies, with a coefficient of 1.35 [95% CI: 1.07 to 1.58]. Among staff, shift 1 and shift 2 personnel were predicted to have higher interaction likelihoods, with coefficients of 1.32 [95% CI: 0.79 to 1.60] and 1.26 [95% CI: 0.58 to 1.64], respectively. Staff on shift 3 had the lowest predicted interaction frequency among staff categories, with a coefficient of 0.74 [95% CI: 0.42 to 1.07]. On the contacting side, residents in low care and medium care were strong predictors of increased contact counts, with coefficients of 1.52 [95% CI: 1.19 to 1.75] and 1.62 [95% CI: 1.28 to 1.91], respectively. Contact initiated by shift 3 staff was associated with a reduced frequency of interactions, reflected in a negative coefficient of –0.38 [95% CI: –0.76 to 0.02].
@@
For the 3-meter contact radius, the GLM analysis averaged across all simulation runs revealed increased interaction likelihoods for low and medium care residents compared to the high care reference group. Specifically, low care residents had a mean coefficient of 1.25 [95% UI: 0.99 to 1.36], while medium care residents showed a higher interaction likelihood with a coefficient of 1.40 [95% UI: 1.09 to 1.52]. Staff on shift 1 exhibited the highest predicted interaction frequency among contactees, with a coefficient of 1.52 [95% UI: 1.06 to 1.69], followed by shift 2 staff with 1.31 [95% UI: 1.10 to 1.48]. Shift 3 staff showed no clear effect, with a coefficient near zero (0.01 [95% UI: -0.60 to 0.94]). When examining interaction frequency from the perspective of the contacter, low care and medium care residents again showed elevated coefficients of 1.65 [95% UI: 1.29 to 1.82] and 1.76 [95% UI: 1.36 to 1.94], respectively. Staff shift 1 and shift 2 also showed moderate increases in contact likelihood as originators of contact (1.22 [95% UI: 0.84 to 1.40] and 1.17 [95% UI: 0.91 to 1.30], respectively). In contrast, shift 3 staff had a negative coefficient of -0.84 [95% UI: -1.35 to -0.13], indicating a substantially lower propensity to initiate contact events.
@@
Figure 3: Predicted contact matrix of different agent type categories (high care resident, low care resident, medium care resident, staff shift 1, staff shift 2 and staff shift 3) per person per day for contact radii of 1.5 metres (A) and 3 metres (B). The matrix was used a generalised linear model with a Poisson distribution to provide an adjusted estimation of contact rates.

### v2

@@ 3.3.4 Activity Generation and Execution
Activity generation for both staff and resident agents is managed through specific functions invoked by the statecharts or other elements of the simulation logic. The activity generation process involves defining both the activation of the specific activity, including the prioritisation and the creation of its token. Tokens appear only in the components of the process model that actively manage an activity’s lifecycle, such as generation, queuing, prioritisation, execution, and therefore do not appear in statechart states or pedestrian movement blocks. Within the agent a process model reads the generation function and locates the activity’s priority. Both staff and resident models utilise priority-based functions to determine the order of task execution. In addition, activity tokens are employed within the model to represent the discrete units of work or actions that an agent must perform. Each token encapsulates the details of a specific activity, including the required subordinate actions, the sequence in which they should occur, location of activity and any associated timing constraints. Tokens serve as structured carriers of activity-specific information, ensuring that agents execute tasks according to predefined parameters in a consistent manner across the population. The management of activity tokens involves queuing, prioritisation, and execution, with tokens queued based on their priority. Each agent’s process model selects the appropriate token for execution, guiding the agent’s actions and ensuring consistency and adherence to the activities’ requirements.
@@ 3.3.5 Social Force Model and Pedestrian Dynamics
A pedestrian simulation using the well characterised Helbing-Molnár-Farkas-Vicsek social force models is used to implement the movement of agents within the simulation environment [16]. This model differs from other social force models by incorporating a more comprehensive set of interaction forces, including individual desires to reach specific destinations while avoiding collisions with other pedestrians and obstacles. It also accounts for the variability in pedestrian behaviour, making it particularly effective for simulating realistic movement patterns in complex environments [16]. The pedestrian model within the simulation is triggered when an activity template requires an agent to navigate to a specific location, such as moving between rooms or transitioning to different areas. The process model within the staff or resident activity generates movement details as part of the task parameters, including the destination, path, and required speed. These parameters are passed to the pedestrian model through the activity templates, which handles the agent’s physical navigation. The pedestrian model ensures realistic movement by managing timing, sequence, and interactions with obstacles and other agents. Using a social force approach, it calculates desired speed, directs movement toward attraction points, and programs agents to follow the shortest routes while avoiding collisions. These features collectively enhance the accuracy of agent interactions and behaviours within the simulation.
@@ 3.3.6 Subcomponent Integration
Once an activity token is selected for execution, the process model coordinates task progression by referencing the associated activity template and invoking the pedestrian model where movement is required. Agent state is updated throughout execution to reflect task progress, interruptions, or completion, ensuring consistency between spatial movement and task logic.
@@ 3.3.7 Feedback Mechanisms
Figure 2: Workflow for task generation, prioritisation, and execution within the agent-based model. Tokens represent individual activity instances as they move through the scheduling process. Blue blocks denote scheduling and control logic, while brown blocks represent physical execution handled by the pedestrian model. Arrows indicate the direction of task flow.
@@ 3.3.8 Integrated Model Operation
For staff, the process models are tailored to manage multiple, often concurrent, activities that may have conflicting priorities. Staff agents are frequently tasked with balancing operational duties, such as logistical support, administrative functions, and direct care activities, in addition to balancing unscheduled events [Table 3]. The process model is designed to adjust execution of tasks as they emerge through priority-based queues, ensuring that tasks are addressed in a logical order based on urgency and importance. Each task is assigned a priority level, and the process model dynamically manages the priority queue, continuously re-evaluating tasks as new ones are generated, or existing tasks are completed. This process model also allows the model to effectively handle conflicting tasks, ensuring the lower priority tasks are either delayed or requeued as necessary. This is particularly important as staff members often juggle multiple operational responsibilities that may require immediate attention. In addition, an importance threshold was assigned in the model to manage the prioritisation-based system. Tasks can be interrupted if they exceed the threshold and are higher priority than the current executing task [Table 3]. If the interrupted task remains within the importance threshold, it is requeued for later completion. The same priority system also contains an Emergency Response category for both staff and residents, which covers rare acute events such as clinical incidents or facility-wide alerts and would supersede all other tasks if triggered. No emergency triggers were scheduled in the scenarios analysed here, so Emergency Response tasks did not occur.
@@ 3.5.1 Identifying Simulated Contacts
We simulated the movement of all the agent types, capturing their interactions and proximity throughout each staff shift. This enabled us to directly compute the instances of close contact among all individuals present, along with the duration of these interactions. In addition, as multiple studies on respiratory illnesses revealed that droplets extend far beyond 1.5 metres which is the recommended social distancing policy for the World Health Organization [5, 4, 39]. As a result, separate contact analyses were performed for contact radii of 1.5 and 3 metres to analyse differences in nonphysical contacts within the space. In the model, a contact is defined as an event where two agents come within 1.5 or 3 metres of each other for a duration equal to or exceeding 3 seconds. Mathematically, a contact between two agents i and j is represented as:
@@
Within the model, a contact is recorded only when two agents are co-located within the same room and within the specified proximity threshold, preventing spurious contacts across physical barriers. Contacts persist across consecutive simulation time steps until the agents separate, after which any subsequent re-encounter is logged as a new contact event. If the same agents come into contact again after a period of separation, a new contact record is created and logged for that pair of agents. This new contact record will have a new start timestamp. The process continues in this manner, with new contact records being generated for each instance of contact between the same pair of agents, and timestamps being updated accordingly. The contact record, once created, includes the following parameters: (i) the identifier of the contacter; (ii) the identifier of the contactee; (iii) the start time of the contact within the effective radius; (iv) the end time of the contact; and (v) the duration of the contact.
@@
$A_{i,k}=\frac{1}{N_{k}}\sum_{j\in\mathcal{G}_{k}}M_{i,j}$
where $i$ denotes a focal individual agent; $k$ denotes an agent type; $\mathcal{G}_{k}$ is the set of agents belonging to type $k$; $N_{k}=|\mathcal{G}_{k}|$ is the number of agents of type $k$; and $M_{i,j}$ represents the total number of contacts between individuals $i$ and $j$.
@@ 3.6 Transmission Risk Simulation
To complement the construction of contact matrices and provide a framework for interpreting their epidemiological implications, a probabilistic airborne transmission model was integrated into the simulation to calculate infection risk based on agent proximity and cumulative contact duration. Transmission probability was dynamically evaluated for each susceptible-infectious agent pair during the simulation based on recorded spatial proximity and accumulated exposure time. Infection outcomes were determined based on the calculated transmission probability at the end of each recorded contact event. Within this model, the transmission component is parameterised for SARS-CoV-2, employing airborne quanta-based transmission estimates and vaccine effectiveness values specific to this pathogen.
@@ 3.6.1 Initial Seeding
To initiate the transmission simulations within the model, an initial infection seeding strategy was employed. For each simulation run, a single staff agent was randomly selected at the beginning of a designated shift to serve as the primary infection case. Only one initial infection was seeded per 24-hour simulation cycle. Separate simulation sets were conducted independently for each shift. Each index case was assumed to be infectious from the start of their respective shift, and all infection events were recorded across the 24-hour simulation period. To allow for direct comparison between scenarios, identical random seeds were used across intervention runs, ensuring that all shared stochastic processes, such as agent movement, task allocation, and seeding, remained consistent, isolating the effect of the intervention. Each simulation run included one infectious staff member as the index case for the 24-hour period.
@@ 3.6.2 Transmission Dynamics
The transmission dynamics presented here are parameterised for SARS-CoV-2, using quanta-generation and inhalation parameters specific to this pathogen. The same formulation can be applied to other airborne diseases by substituting the corresponding pathogen-specific parameters. A probabilistic transmission risk model, applying a proximity-based Wells-Riley formulation, was integrated into the simulation [32]. Transmission risk between agents was calculated based on direct proximity and cumulative contact duration, without assuming homogeneous aerosol mixing or dilution effects through room-scale ventilation. The transmission probability $P_{i,j}$ from an infectious individual i to a susceptible individual j was defined as:
@@
Where $\lambda$ is the effective transmission rate per second of close contact, calculated as the product of the quanta generation rate (I) and the inhalation rate (q) and normalised to a per-second basis; $t_{i,j}$ represents the cumulative contact duration, measured in seconds, between individuals i and j, derived directly from simulation-recorded interaction data [Table 5]
@@
I Quanta generation rate 25 quanta/hr (95% CI: 14–54) [6, 24, 10]
q Inhalation rate of susceptible individuals during sedentary activities 0.5 m3/hr (95% CI: 0.48–0.54) [6, 15, 11]
@@ 3.6.4 Vaccination Intervention Scenarios
Given the elevated vulnerability of aged care settings and the recognised role of booster vaccination in mitigating infection risk, intervention scenarios incorporating resident and staff booster coverage were evaluated. Vaccination status was incorporated into the simulation as an agent attribute assigned at initialisation. For each simulation run, staff agents were assigned a binary vaccination status according to the intervention scenario being evaluated. Vaccination status remained fixed throughout the 24-hour simulation cycle. In this study, the baseline condition (“no booster administered”) represents a population whose immunity from the primary vaccination series has waned and is therefore modelled as having no effective protection at the time of exposure.
During the evaluation of transmission probabilities, vaccination status was incorporated as a scaling adjustment applied to the calculated probability of infection for susceptible agents. If a susceptible agent was recorded as vaccinated, their computed transmission probability $P_{\text{transmission}}(i,j)$ was adjusted as:
@@
Booster vaccination was modelled assuming receipt of the Pfizer-BioNTech Comirnaty Omicron XBB.1.5 formulation. This vaccine was selected as it represented the predominant COVID-19 booster administered in Australia during the study period, consistent with Australian Technical Advisory Group on Immunisation recommendations prioritising mRNA-based vaccines for booster doses and confirmed in national policy documents and vaccine rollout reports [2, 1, 35]. Vaccine effectiveness against any SARS-CoV-2 infection for the Comirnaty Omicron XBB.1.5 booster was assumed to be 70.6% (95% CI: 43.3%–84.8%), based on recently reported effectiveness estimates encompassing both symptomatic and asymptomatic infections [37, 34, 7]. Vaccination status modified the probability of infection at the time of contact evaluation without altering other agent attributes, including mobility patterns, task execution, or contact behaviours.
Within the simulation, vaccination status modified transmission probability only at the point of contact event evaluation. At the conclusion of each cumulative contact between an infectious and susceptible agent, the adjusted transmission probability $P_{\text{transmission}}^{*}(i,j)$ was computed based on the susceptible agent’s vaccination status. A uniform random number $U\sim\mathcal{U}(0,1)$ was then drawn to determine infection status, following the method described above. Three vaccination intervention scenarios were evaluated: (i) no booster vaccination among staff or residents (baseline); (ii) booster vaccination administered to all residents but not staff; and (iii) booster vaccination administered to both staff and residents. Vaccination status was assigned at model initialisation and fixed throughout the 24-hour simulation cycle. All infection probability adjustments were applied based on vaccination status as previously described.
@@ 4.2 Contact Matrix
For the contact matrix generation, the simulation ran for 5 days, starting at 6:15 AM on day 1 and ending at 6:15 AM on day 6. During this period, multiple data series were collected for the analysis, specifically focusing on two different contact radii: 1.5 meters and 3 meters [Figure 3]. At a 1.5-meter contact radius, simulated interactions among residents showed variation across different care levels. Residents in low care had the highest predicted intra-group interactions, averaging 64.4 contacts [95% UI: 52.5–80.1], followed closely by medium care residents with 49.6 contacts [95% UI: 41.7–56.3]. High care residents had substantially fewer intra-group contacts, averaging 3.5 [95% UI: 2.3–5.0]. Staff-resident interactions were most frequent for shift 1 and shift 2 care workers, particularly with medium care residents (60.9 [95% UI: 45.0–73.3] and 58.1 [95% UI: 33.3–72.0] contacts, respectively). Contacts with low care residents were also high for shift 1 (54.3 [95% UI: 41.3–61.9]) and shift 2 (52.0 [95% UI: 28.2–64.6]). In contrast, shift 3 staff demonstrated the lowest levels of interaction across all resident types, with resident contacts not exceeding 35 contacts across any group and intra-shift contacts averaging only 2.5 [95% UI: 0.6–6.4].
@@
At a 3-meter contact radius, the simulated interaction patterns largely mirrored those observed at 1.5 meters, though overall contact frequencies were slightly increased in most groups. Low care residents had the highest intra-group interactions, averaging 64.8 contacts [95% UI: 45.3–69.1], followed closely by medium care residents with 51.0 contacts [95% UI: 47.8–52.1]. High care residents continued to exhibit fewer intra-group contacts, averaging 3.0 [95% UI: 2.5–4.6]. Among staff, shift 1 had the most frequent predicted interactions, particularly with medium care residents (72.9 [95% UI: 49.3–81.2]) and low care residents (66.1 [95% UI: 44.8–72.1]). Shift 2 staff showed similar patterns, with 60.0 contacts with medium care residents [95% UI: 49.9–65.2] and 54.4 with low care residents [95% UI: 52.5–59.1]. Shift 3 staff maintained the lowest levels of interaction across all groups, with resident contacts not exceeding 36 contacts and intra-shift contacts averaging only 4.7 [95% UI: 4.2–6.0].
@@
For the 1.5-meter contact radius, the GLM analysis applied to the simulated contact data revealed that low care residents had a significantly greater likelihood of being involved in interactions compared to high care residents, with a coefficient of 1.35 [95% CI: 1.07 to 1.58. Medium care residents also showed elevated predicted interaction frequencies, with a coefficient of 1.21 [95% CI: 0.97 to 1.37]. Among staff, shift 1 and shift 2 personnel were predicted to have higher interaction likelihoods, with coefficients of 1.32 [95% CI: 0.79 to 1.60] and 1.26 [95% CI: 0.58 to 1.64], respectively. Staff on shift 3 had the lowest predicted interaction frequency among staff categories, with a coefficient of 0.74 [95% CI: 0.42 to 1.07]. On the contacting side, residents in low care and medium care were strong predictors of increased contact counts, with coefficients of 1.52 [95% CI: 1.19 to 1.75] and 1.62 [95% CI: 1.28 to 1.91], respectively. Contact initiated by shift 3 staff was associated with a reduced frequency of interactions, reflected in a negative coefficient of –0.38 [95% CI: –0.76 to 0.02].
@@
For the 3-meter contact radius, the GLM analysis averaged across all simulation runs revealed increased interaction likelihoods for low and medium care residents compared to the high care reference group. Specifically, low care residents had a mean coefficient of 1.40 [95% UI: 1.09 to 1.52] , while medium care residents showed a higher interaction likelihood with a coefficient of 1.25 [95% UI: 0.99 to 1.36]. Staff on shift 1 exhibited the highest predicted interaction frequency among contactees, with a coefficient of 1.52 [95% UI: 1.06 to 1.69], followed by shift 2 staff with 1.31 [95% UI: 1.10 to 1.48]. Shift 3 staff showed no clear effect, with a coefficient near zero (0.01 [95% UI: -0.60 to 0.94]). When examining interaction frequency from the perspective of the contacter, low care and medium care residents again showed elevated coefficients of 1.65 [95% UI: 1.29 to 1.82] and 1.76 [95% UI: 1.36 to 1.94], respectively. Staff shift 1 and shift 2 also showed moderate increases in contact likelihood as originators of contact (1.22 [95% UI: 0.84 to 1.40] and 1.17 [95% UI: 0.91 to 1.30], respectively). In contrast, shift 3 staff had a negative coefficient of -0.84 [95% UI: -1.35 to -0.13], indicating a substantially lower propensity to initiate contact events.
@@
Figure 3: Predicted contact matrix showing mean contacts per person per day between agent type categories: high care residents, medium care residents, low care residents, and staff on shifts 1–3, for contact radii of 1.5 metres (A) and 3 metres (B). Contact rates were estimated using a generalised linear model with a Poisson distribution to adjust for differences in group size and stochastic variability.